# Voting by mail: a Markov chain model for managing the security risks of absentee voting

### v2

@@ Abstract
The scrutiny surrounding vote-by-mail (VBM) in the United States has increased in recent years, raising concerns about the integrity and security of absentee voting. This paper addresses these issues by introducing a dynamic mathematical modeling framework for performing a risk assessment of VBM processes. We introduce a discrete-time Markov chain (DTMC) to model the VBM process and assess election performance and risk with a novel layered network approach that considers the interplay between VBM processes, malicious and non-malicious threats, and security mitigations. The time-inhomogeneous DTMC framework captures dynamic risks and evaluates performance over time. The DTMC model accounts for a spectrum of outcomes, from unintended voter errors to sophisticated, targeted attacks, representing a significant advancement in the risk assessment of VBM planning and protection. A case study based on real-world data from Milwaukee County, Wisconsin, is used to evaluate the DTMC model. The analysis includes the development of attack scenarios and the evaluation of security measures, to illustrate the impact of different attack timings. The analysis suggests that ballot drop boxes and automatic ballot notification systems are crucial for ensuring secure and reliable operations.
@@ 1 Introduction
The popularity of VBM among voters has highlighted the need for an efficient, reliable, and secure plan for this voting modality. The increased volume of VBM ballots in the 2020 General Election revealed issues such as mailing delays, security concerns, and accessibility to voting infrastructure. Since 2020, Colorado, Oregon, and six other states have adopted “all-mail” voting, which involves pro-actively sending mail-in ballots to all registered voters [18]. Additionally, many US states have proposed or are considering legislation that has implications for VBM operations and security [17]. Election planning requires a moderate understanding of the security needs crucial for VBM processes and, by extension, core election infrastructure.
@@
The VBM process involves a complex interplay of time-sensitive tasks among election offices, voters, and the USPS. These tasks comprise the election procedures and policies managed at the state and local levels. As a result, VBM processes can vary depending on the location. There is no practical way to physically test absentee election operations at scale before an election, which drives the need for enhanced security planning tools and risk analysis models. Additionally, the VBM process occurs over several months, making it vulnerable to temporal issues occurring at specific times, such as USPS mail processing delays causing ballots to arrive after an election day when they are no longer counted (in many US states), as well as in-person absentee voting that is only available on specific days leading up to an election. We integrate these risks into a dynamic mathematical model of the VBM system capable of evaluating different threats and measures to mitigate them.
@@
This paper presents a methodology to support VBM risk analysis by introducing a novel discrete-time Markov chain (DTMC) for VBM to capture dynamic election performance. The modeling approach uses the attack tree models of malicious (intentional) and non-malicious (unintentional) risks and security mitigations. Attack tree models have been used to characterize risks in prior studies on voting by mail security [34, 24]. A novel aspect of the DTMC framework is that it models the linkages between attack trees, security mitigations, and VBM processes to characterize and predict dynamic VBM performance under various threat scenarios and operating conditions. Furthermore, the DTMC model leverages attack trees that provide a structured approach to capturing malicious and non-malicious risks to VBM systems. We focus on the number of ballots affected by election risks, not on the likelihood of the outcome of an election changing, since any inaccuracies weaken the legitimacy of a political system, regardless of whether these inaccuracies change election outcomes. The DTMC framework highlights how different types, intentions, and timings of attacks impact overall VBM performance and security. We acknowledge that successful attacks could change the outcome of an election. Using real-world data from Milwaukee County in the 2020 General Election, we evaluate VBM performance under different conditions and policy scenarios. Contributions of this paper include the following.
@@
We formulate a new DTMC using a layered network approach that evaluates election performance and risk over time.
@@
We perform a detailed analysis of the DTMC model using a real-world case study based on Milwaukee County, Wisconsin. We study multiple threat scenarios to shed light on how mitigation strength impacts performance of the VBM process.
@@ 2 Literature review
Research on the allocation of voting machines to polling locations for in-person voting has been extensively explored through various methodologies. One notable example is a study by Yang et al. [41] who uses discrete event simulation to model election day processes, combining queuing theory and optimization to address resource concerns. Similarly, Li et al. [11] employ simulation optimization to assess the impact of voting machine distribution on voter queue times. The significance of these studies extends to a tailored risk assessment for VBM. Schmidt and Albert [25] apply discrete event simulation to understand the effects of safety measures implemented during the COVID-19 pandemic on wait times and other election performance indicators. In addition, McIntyre [12] focus on optimizing the polling locations to reduce voter waiting times. These approaches underscore the importance of modeling and optimization in improving election processes, informing the development of risk assessment methodologies for election infrastructure. Specific to the VBM system, Scala et al. [23] develop a detailed model of the VBM process, which meticulously outlines its various components and the physical journey of ballots. The Cybersecurity and Infrastructure Security Agency (CISA) conducted a comprehensive assessment of mail-voting security for the 2020 General Election [5], contributing to efforts to identify and mitigate VBM-related risks during the pandemic.
@@
A growing body of literature characterizes threats to the voting process using attack trees. Attack trees help identify and prioritize potential vulnerabilities in a system, making them an essential tool in cybersecurity planning and risk mitigation. Attack trees are fundamental in cybersecurity for visualizing vulnerabilities and potential attacker paths [27]. An attack tree starts with the primary goal of the attacker (the root) and branches out into different methods or steps that the attacker might employ (the leaves). This branching structure often incorporates logic gates like AND and OR to demonstrate how different actions might combine to lead to the attack goal. Terminal attack nodes, are the leaf nodes of an attack tree that initiate the attack on the system. The University of South Alabama [34] present a risk analysis report for election systems and models security threats to the VBM using attack trees that capture all possible vulnerabilities in the VBM system. Scala et al. [24] expand on the VBM threat trees and adds to tree logic enumerating all terminal attack nodes to account for the new mechanisms available, such as drop boxes and in-person absentee voting. Haseltine et al. [9] build upon this body of literature to identify mitigation strategies and policies that protect the VBM process, accounting for linkages to attack trees relevant to VBM processes. Our DTMC framework utilizes these attack trees and mitigations.
@@ 3.1 Process Layer
First, we introduce seven process states that capture the movement of ballots in the process layer and define the operation of the VBM system. Each of these process states is also a state in the DTMC, and the states represent the change in physical location of the ballot over time. The voter requests a ballot in state , the election office then fulfills this request in state and mails an unmarked ballot to the voter. In state the unmarked ballot is handled by the United States Postal Service (USPS) and is in transit to the voter. In state the unmarked ballot reaches the voter, however, it is up to the voter when the marked ballot is filled out and returned. In state the voter selects the method of return for the marked ballot. Standard operation for returning a ballot is a transition to state via the USPS. Alternatively, the voter could return the ballot by drop box, if drop boxes are available. The final process state occurs when the marked ballot reaches the election office and is held until processing on election day. The ballot process is summarized by the following states:
@@ 3.2 Final Ballot States
At the end of the VBM process, ballots move to a final absorbing state that reflects the ballot status and is used to evaluate election performance. The final states capture all possible final states the voter ballot can take on at the end of the process, and they are defined to be mutually exclusive. The voter ballots can be counted (C) or not counted (NC) by the election officials. There are several reasons a ballot may not be counted, and not all result from attacks to the process. Some ballots are not counted simply because they arrive at the election office after the required date for processing (L), which is typically an election day.
@@
(NC,L): Voter ballot is returned and does not arrive at the election office by election day, designated late ballot.
(C,A): Voter ballot is maliciously altered, accepted, and counted on election day
(NC,A): Voter ballot is maliciously altered, however it is rejected and not counted on election day.
(NC,NR): Voter does not return the ballot, includes ballots lost in transit and ballots never delivered [40].
@@ 3.3 Attacks Layer
The attacks layer is comprised of active vulnerabilities to the VBM process and can reflect different attacker goals, including changing the outcome of an election or eroding trust in political systems. Vulnerabilities in the VBM system are documented in attack trees by the US Election Assistance Commission [34] and have been expanded upon since the 2020 General Election [23]. The terminal nodes of these attack trees represent “access” points to additional vulnerabilities within the process. For this reason, we model each of these terminal leaf nodes of the attack trees. Attack trees, a valuable tool in cybersecurity, help assess threats to a system [27]. They depict potential vulnerabilities and attacker paths in a Boolean logic tree structure, where the root node represents the ultimate goal of an attacker. Attack trees use a combination of Boolean AND() and OR() logic gates to trace paths from terminal attack nodes back to the root node. Figure 1 shows an example of an attack tree that visualizes two attacks corresponding to the malicious loss () and accidental loss () of a ballot. In Figure 1, terminal attack nodes and form part of an attack path leading to a successful insider attack, represented by root node 1 with only OR-gates along the paths. These attack paths offer a static view of threats within the VBM process, with the terminal leaf nodes and serving as entry points.
@@
Haseltine et al. [9] explore the impact of attacks on the VBM process, focusing on how each attack could affect voter ballots in terms of being lost, late, or maliciously altered. Expanding upon this, here we classify attacks not only based on their impact on ballots but also on the VBM process as a whole. Terminal attack nodes are classified into three categories: fail rate increase, where attacks change the probability associated with an undesirable outcome; ballot altering, which involves direct tampering with the voter ballot; and process-altering, where attacks cause voter ballots to deviate from the standard physical VBM process. This broader perspective allows for a more comprehensive understanding of the attacks’ implications.
@@
Any attack that affects the probability associated with an outcome, e.g., a transition probability in the DTMC model, is considered to be a fail-rate increase. These attacks can be non-malicious, e.g., caused by voter errors when filling out a ballot. As a result, the election office has a higher rejection rate for returned ballots, resulting in uncounted votes. However, these attacks can also be malicious, e.g., caused by a bad actor at the election office erroneously failing ballots. For example, consider a malicious insider attack () in which a VBM ballot is erroneously deemed insufficiently filled out (errant failed signature). If the malicious attack occurs, it increases the probability that a completed ballot is rejected.
@@
Changing the vote on a marked ballot alters the information in the original ballot we consider a “ballot-altering attack.” Masquerade attacks involve a bad actor maliciously altering a ballot. These attacks are more complex in that they create alternate ballot states in the DTMC that are needed to reflect the ballot’s change in composition. For example, a masquerade attack in which a bad actor votes on behalf of someone in a central housing situation requires several steps. The malicious actor must first register to vote on behalf of the original voter, intercept their mail, mark the ballot, and return the altered ballot to the VBM process. We designated this type of attack as ballot modifying. It is important to note that modified ballots still have the potential to be counted. For this reason, there is an alternate path for the modified ballots in the DTMC. This is later illustrated in the DTMC model attack layer in Section 4, which reflects the VBM process states in the process layer to indicate that the ballots of those states were maliciously altered during transit.
@@
In the VBM system, there are seven distinct states that an original ballot must pass through to be counted and remain unaltered (see Section 3.1). When someone attempts to disrupt this process, it is known as a process-altering attack. For example, suppose an election volunteer steals the marked ballot while it is in transit to the election office. This malicious attack alters the ballot process by holding the ballot in an “attack state” outside of the normal stages of the election process until the voter is notified and uses a recourse action such as requesting a replacement ballot, or it is deemed not returned and is not counted. The notable implication of process-altering attacks is that they can delay or prevent ballots from reaching the election office.
@@
Table 1 presents the list of terminal attack nodes to the VBM system from Scala et al. [23]. We focus on those attacks that require voter or election office recourse actions to be countered. The first column lists the updated terminal attack nodes in the VBM attack tree, while the second column identifies the specific stage in the VBM process affected by each attack. The third column indicates the final state of the ballot if the attack is successful. Table 1 delineates the intent and classification of each terminal node attack in the last two columns, offering insights into the nature of these threats. Table 1 boldface rows are considered the most significant risks to VBM. These attacks are considered significant attacks, since they introduce substantial variations in the final ballot states, thereby highlighting the relevance to our analysis.
@@
There are malicious and non-malicious attack types. Malicious attacks are targeted and have a high local impact for a limited period of time (e.g., a day) of being active. Conversely, non-malicious attacks (e.g., voter error and accidental loss) are accidental and could occur any time in the DTMC model time horizon. For example, in the 2020 General Election, some ballots were not counted due to voters failing to sign or bundle the ballots correctly (attack ). This non-malicious attack results in ballots being rejected by the election office, thus inhibiting the VBM process if not countered. We express this attack as a small probability of occurrence for each ballot over the entire evaluation period.

### v3

@@ Abstract
The scrutiny surrounding vote-by-mail (VBM) in the United States has increased in recent years, highlighting the need for a rigorous quantitative framework to evaluate the resilience of the absentee voting infrastructure. This paper addresses these issues by introducing a dynamic mathematical modeling framework for performing a risk assessment of VBM processes. We introduce a discrete-time Markov chain (DTMC) to model the VBM process and assess election performance and risk with a novel layered network approach that considers the interplay between VBM processes, malicious and non-malicious threats, and security mitigations. The time-inhomogeneous DTMC framework captures dynamic risks and evaluates performance over time. The DTMC model accounts for a spectrum of outcomes, from unintended voter errors to sophisticated, targeted attacks, representing a significant advancement in the risk assessment of VBM planning and protection. A case study based on real-world data from Milwaukee County, Wisconsin, is used to evaluate the DTMC model. The analysis includes hypothetical worst-case attack scenarios to stress-test VBM processes and to assess the efficacy of security measures and the impact of different attack timings. The analysis suggests that ballot drop boxes and automatic ballot notification systems are crucial for reducing the attack surface to ensure secure and reliable operations.
@@ 1 Introduction
The popularity of VBM among voters has highlighted the need for an efficient, reliable, and secure plan for this voting modality. The increased volume of VBM ballots in the 2020 General Election revealed issues such as mailing delays, security concerns, and accessibility to voting infrastructure. Since 2020, Colorado, Oregon, and six other states have adopted “all-mail” voting, which involves pro-actively sending mail-in ballots to all registered voters [20]. Additionally, many US states have proposed or are considering legislation that has implications for VBM operations and security [19]. While election processes are believed to be secure in the United States [12, 38], election planning requires an understanding of the security needs crucial for VBM processes and, by extension, core election infrastructure.
@@
The VBM process involves a complex interplay of time-sensitive tasks among election offices, voters, and the USPS. These tasks comprise the election procedures and policies managed at the state and local levels. As a result, VBM processes can vary depending on the location. While there is extensive preparation prior to every election, it is not practical to physically stress-test the entire VBM infrastructure, which drives the need for enhanced security planning tools and risk analysis models. Additionally, the VBM process occurs over several months, making it vulnerable to temporal issues occurring at specific times, such as USPS mail processing delays causing ballots to arrive after an election day when they are no longer counted (in many US states), as well as in-person absentee voting that is only available on specific days leading up to an election. We integrate these risks into a dynamic mathematical model of the VBM system capable of evaluating different threats and measures to mitigate them.
@@
This paper presents a methodology to support VBM risk analysis by introducing a novel discrete-time Markov chain (DTMC) for VBM to capture dynamic election performance and stress-test VBM processes. The modeling approach uses the attack tree models of malicious (intentional) and non-malicious (unintentional) risks and security mitigations. Attack tree models have been used to characterize risks in prior studies on voting by mail security [39, 25]. A novel aspect of the DTMC framework is that it models the linkages between attack trees, security mitigations, and VBM processes to characterize and predict dynamic VBM performance under various threat scenarios and operating conditions.
The framework captures a predefined set of threats and VBM process rules to examine hypothetical worst-case scenarios to test the ability of security mitigations to withstand these targeted attacks independent of the historical rarity of such events. The case study analyses seek to stress-test the current security landscape rather than provide a forecast of daily operations. The DTMC framework highlights how different types, intentions, and timings of attacks impact overall VBM performance and security. We acknowledge that successful attacks could potentially change the outcome of an election. Using real-world data from Milwaukee County in the 2020 General Election, we evaluate VBM performance under different conditions and policy scenarios.
We focus on the number of ballots affected by election risks, not on the likelihood of the outcome of an election changing, since any inaccuracies weaken the legitimacy of a political system, regardless of whether these inaccuracies change election outcomes. From a decision analysis perspective, minimizing the expected number of ballots affected by risk serves as a means objective. While the primary fundamental objective of an adversary is typically viewed as preserving the election outcome (i.e., minimizing the probability that the true winner is not certified), other fundamental objectives for an adversary may be to undermine public trust in the democratic process or delegitimize the results, regardless of the final count [33, 37]. Under the trust erosion objective, for example, the adversary does not need to mathematically flip the election outcome; they only need to generate enough affected ballots (e.g., lost, stolen, or successfully challenged votes) to fuel a narrative of systemic failure. Therefore, our focus on the number of affected ballots supports a robust defense strategy, since it simultaneously hardens the system against outcome-reversal attacks, which require affected volume to exceed the margin of victory, and trust-erosion attacks, which require affected volume to exceed the public’s tolerance for error.
Contributions of this paper include the following.
@@
We formulate a new DTMC using a layered network approach that evaluates election performance and risk over time. The model is derived from an existing VBM process map [24], which aggregates granular operational steps into functional states to allow for a quantification of ballot flow and risk measurement across the voting timeline.
@@
We introduce a methodology for stress-testing VBM processes. Given that historical security incidents such as confirmed fraud cases are statistically rare, empirical threat probabilities are difficult to estimate. Therefore, we consider hypothetical worst-case scenarios to evaluate the system’s resilience against potential disruptions.
4.
We perform a detailed analysis of the DTMC model using a real-world case study based on Milwaukee County, Wisconsin that considers multiple, hypothetical threat scenarios to stress-test the model and shed light on how mitigations impact performance of the VBM process.
@@ 2 Literature review
Crimmins et al.[5] use robust optimization to develop efficient logic and accuracy testing to ensure that voting machines accurately count votes.
@@
Research on the allocation of voting machines to polling locations for in-person voting has been extensively explored through various methodologies. One notable example is a study by Yang et al. [46] who uses discrete event simulation to model election day processes, combining queuing theory and optimization to address resource concerns. Similarly, Li et al. [13] employ simulation optimization to assess the impact of voting machine distribution on voter queue times. The significance of these studies extends to a tailored risk assessment for VBM. Schmidt and Albert [26] apply discrete event simulation to understand the effects of safety measures implemented during the COVID-19 pandemic on wait times and other election performance indicators. In addition, McIntyre [14] focus on optimizing the polling locations to reduce voter waiting times. These approaches underscore the importance of modeling and optimization in improving election processes, informing the development of risk assessment methodologies for election infrastructure.
Many ballots by mail are tallied by scanning ballots by voting equipment. To ensure that votes are accurately counted, Crimmins et al. [5] use robust optimization to develop efficient logic and accuracy testing. Specific to the VBM system, Scala et al. [24] develop a detailed model of the VBM process, which meticulously outlines its various components and the physical journey of ballots. The Cybersecurity and Infrastructure Security Agency (CISA) conducted a comprehensive assessment of mail-voting security for the 2020 General Election [36], contributing to efforts to identify and mitigate VBM-related risks during the pandemic.
@@
A growing body of literature characterizes threats to the voting process using attack trees. Attack trees help identify and prioritize potential vulnerabilities in a system, making them an essential tool in cybersecurity planning and risk mitigation. Attack trees are fundamental in cybersecurity for visualizing vulnerabilities and potential attacker paths [28]. An attack tree starts with the primary goal of the attacker (the root) and branches out into different methods or steps that the attacker might employ (the leaves). This branching structure often incorporates logic gates like AND and OR to demonstrate how different actions might combine to lead to the attack goal. Terminal attack nodes, are the leaf nodes of an attack tree that initiate the attack on the system. The U.S. Election Assistance Commission [39] present a risk analysis report for election systems and models security threats to the VBM using attack trees that capture all possible vulnerabilities in the VBM system. Scala et al. [25] expand on the VBM threat trees and adds to tree logic enumerating all terminal attack nodes to account for the new mechanisms available, such as drop boxes and in-person absentee voting. Haseltine et al. [9] build upon this body of literature to identify mitigation strategies and policies that protect the VBM process, accounting for linkages to attack trees relevant to VBM processes. Our DTMC framework utilizes these attack trees and mitigations.
@@ 3.1 Process Layer
First, we introduce seven process states that capture the movement of ballots in the process layer and define the operation of the VBM system. Each of these process states is also a state in the DTMC, and the states represent the change in physical location of the ballot over time. The voter requests a ballot in state , the election office then fulfills this request in state and mails an unmarked ballot to the voter. We implicitly assume that all ballots are sent by request to voters by mail, although in some US states, a voter may pick up an absentee ballot in person at the Board of Elections and return it later. In state the unmarked ballot is handled by the United States Postal Service (USPS) and is in transit to the voter. In state the unmarked ballot reaches the voter, however, it is up to the voter when the marked ballot is filled out and returned. In state the voter selects the method of return for the marked ballot. Standard operation for returning a ballot is a transition to state via the USPS. Alternatively, the voter could return the ballot by drop box, if drop boxes are available, which allows the associated ballot to bypass stage VI. The final process state occurs when the marked ballot reaches the election office and is held until processing on election day. The ballot process is summarized by the following states:
@@ 3.2 Final Ballot States
At the end of the VBM process, ballots move to a final absorbing state that reflects the ballot status and is used to evaluate election performance. The final states capture all possible final states the voter ballot can take on at the end of the process, and they are defined to be mutually exclusive. The voter ballots can be counted (C) or not counted (NC) by the election officials. There are several reasons a ballot may not be counted, and not all result from attacks to the process. Some ballots are not counted simply because they arrive at the election office after the required date for processing (L), which is typically an election day. Ballots are not counted if they are missing the voter’s signature, the signature does not match the signature on file, or the ballot is not properly sealed in the envelope or secrecy sleeve.
@@
(NC,L): Voter ballot is returned and either fails to arrive by the election day or is postmarked after the election day (per relevant voting law).
(C,A): Voter ballot is maliciously altered, accepted, and counted.
(NC,A): Voter ballot is maliciously altered, and it is rejected and not counted.
(NC,NR): Voter does not return the ballot, includes ballots lost in transit and ballots never delivered [45], which typically triggers the provisional voting process.
@@ 3.3 Attacks Layer
The attacks layer is comprised of active vulnerabilities to the VBM process and can reflect different attacker goals, including changing the outcome of an election or eroding trust in political systems. Vulnerabilities in the VBM system are documented in attack trees by the US Election Assistance Commission [39] and have been expanded upon since the 2020 General Election [24]. The terminal nodes of these attack trees represent “access” points to additional vulnerabilities within the process. For this reason, we model each of these terminal leaf nodes of the attack trees and they are labeled numerically with a prefix of “X,” using the notation from Scala et al. [24]. For example represents the terminal attack node associated with the malicious loss of a ballot. Attack trees, a valuable tool in cybersecurity, help assess threats to a system [28]. They depict potential vulnerabilities and attacker paths in a Boolean logic tree structure, where the root node represents the ultimate goal of an attacker. Attack trees use a combination of Boolean AND() and OR() logic gates to trace paths from terminal attack nodes back to the root node. Figure 1 shows an example of an attack tree that visualizes two attacks corresponding to the malicious loss () and accidental loss () of a ballot [39]. In Figure 1, terminal attack nodes and form part of an attack path leading to a successful insider attack, represented by root node 1 with only OR-gates along the paths. These attack paths offer a static view of threats within the VBM process, with the terminal leaf nodes and serving as entry points.
@@
Haseltine et al. [9] explore the impact of attacks on the VBM process, focusing on how each attack could affect voter ballots in terms of being lost, late, or maliciously altered. Expanding upon this, here we classify attacks not only based on their impact on ballots but also on the VBM process as a whole. Terminal attack nodes are classified into three categories: fail rate increase, where attacks change the relative likelihood of an undesirable outcome; ballot altering, which involves direct tampering with the voter ballot; and process-altering, where attacks cause voter ballots to deviate from the standard physical VBM process. This broader perspective allows for a more comprehensive understanding of the attacks’ implications.
@@
Any attack that affects the relative likelihood an outcome, e.g., a transition in the DTMC model, is considered to be a fail-rate increase. These attacks can be non-malicious, e.g., caused by voter errors when filling out a ballot. As a result, the election office has a higher rejection rate for returned ballots, resulting in uncounted votes. However, these attacks can also be malicious, e.g., caused by a bad actor at the election office erroneously failing ballots. For example, consider a malicious insider attack () in which a VBM ballot is erroneously deemed insufficiently filled out (errant failed signature). If the malicious attack occurs, it increases the relative likelihood that a completed ballot is rejected.
@@
Changing a marked ballot alters the information in the original ballot, which we consider a “ballot-altering attack.” Masquerade attacks involve a bad actor maliciously altering a ballot. These attacks are more complex in that they create alternate ballot states in the DTMC that are needed to reflect the ballot’s change in composition. For example, a masquerade attack in which a bad actor votes on behalf of someone in a central housing situation requires several steps. In this type of attack, which is believed to be extremely rare, the malicious actor must first register to vote on behalf of the original voter, intercept their mail, mark the ballot, and return the altered ballot to the VBM process. We designated this type of attack as ballot modifying. It is important to note that modified ballots still have the potential to be counted. For this reason, there is an alternate path for the modified ballots in the DTMC. This is later illustrated in the DTMC model attack layer in Section 4, which reflects the VBM process states in the process layer to indicate that the ballots of those states were maliciously altered during transit.
@@
In the VBM system, there are seven distinct states that an original ballot must pass through to be counted and remain unaltered (see Section 3.1). When someone attempts to disrupt this process, it is known as a process-altering attack. For example, for illustration purposes suppose an election volunteer steals the marked ballot while it is in transit to the election office. This unlikely malicious attack would alter the ballot process by holding the ballot in an “attack state” outside of the normal stages of the election process until the voter is notified and uses a recourse action such as requesting a replacement ballot, or it is deemed not returned and is not counted. The notable implication of process-altering attacks is that they can delay or prevent ballots from reaching the election office.
@@
Table 1 lists the terminal attack nodes to the VBM system from Scala et al. [24] that require voter or election office recourse actions to be countered. The first column lists the updated terminal attack nodes in the VBM attack tree, while the second column identifies the specific stage in the VBM process affected by each attack. The third column indicates the final state of the ballot if the attack is successful. Table 1 delineates the intent and classification of each terminal node attack in the last two columns, offering insights into the nature of these threats. The terminal attack notes in boldface in Table 1 are considered to represent the most significant risks to VBM according to a detailed risk analysis by Scala et al. [25], and therefore, we model these terminal attack nodes in our approach. The model can be extended in a straightforward manner to include other terminal attack nodes.
@@
There are malicious and non-malicious attack types. Malicious attacks are targeted and have a high local impact for a limited period of time (e.g., a day) of being active. Conversely, non-malicious attacks (e.g., voter error and accidental loss) are accidental and could occur any time over the DTMC model time horizon. For example, in the 2020 General Election, some ballots were not counted due to voters failing to sign or bundle the ballots correctly (attack ). This non-malicious attack results in ballots being rejected by the election office, thus inhibiting the VBM process if not countered. We express this attack as a low-likelihood occurrence for each ballot over the entire evaluation period.